1.00 mL of a 3.80 × 10–4 M solution of oleic acid is diluted with 9.00 mL of petroleum ether, forming solution A. Then 2.00 mL of solution A is diluted with 8.00 mL of petroleum ether, forming solution B. What is the concentration of solution B?
1
Expert's answer
2019-11-01T05:44:54-0400
"M_1
V_1 = M_2
V_2 \\ M_2 =\frac{ M_1
V_1}
{V_2}"

After first dilution,


"M_2=\frac {(3.95\times10^{−4 }mol\ 
L^{-1} )(0.001L)}
{.0100L} = 3.95\times10^{−5} M"

After second dilution,

"M_3=\frac{(3.95\times10^{-5})(0.002)}{0.010}=7.90\times10^{-6}\ M"